DQuestion 5 3 pts A firm with yearly net income of $33,000 has only two customers, each of whom makes identical purchases once per year. Timely Tom makes his purchase today and at the start of next year, while Late Louie makes his purchase half a year from today and halfway into next year (18 months from today). Moreover, Late Louie always pays 6 months late, i.e. at year end. What is the NPV of all cash flows to this firm in the next two years? Assume a yearly discount rate of 10%. D Question 6 3 pts Consider the following debt-free project for a machine: Physical Lifespan Capital Expense, Year O EBITDA/ Year Cost of Capital Tax Rate Depreciation Life 7 years $14,000 $8,000 9.9% 39% 4 years What is the value of this project? Please answer both

Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

Answer:

Question 5)

Yearly income =33000

Since two consumer makes the identical purchase

So income from Timley tom=33000/2=16500

So income from Late louie =33000/2=16500

Discount rate r=10%

Cash flow diagram for firm is

Year Timely Tom T Late louie L Total P=T+L
0 16500 0 16500
1 16500 16500 (since louie is paying at end of year) 33000
2 0 16500 16500

NPV=sum of PV all Cash flows

PV=P/(1+r)^n

n=year in which cash flow occur

NPV=16500/(1+10%)^0+33000/(1+10%)^1+16500/(1+10%)^2=182863.64